OpenStudy (anonymous):

the perimeter of a semi-circle is 90cm.Calculate its area @ganeshie8 @girlover

OpenStudy (anonymous):

that's wrong @botype

OpenStudy (anonymous):

according to the book it is 481.25

OpenStudy (anonymous):

but how?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@sleepyjess @Directrix

OpenStudy (sleepyjess):

ummmmmmmmmmm.......... @pooja195

pooja195 (pooja195):

if u had a pic it would be easier to help

OpenStudy (anonymous):

only the wording was given

Directrix (directrix):

The perimeter or circumference of a circle is given by C = 2 * pi * r You have half a perimeter in the semicircle so 90 = pi * r

Directrix (directrix):

Solve this for r: 90 = pi * r Then, you can use the area of a circle formula to get the area of the semicircle assuming the semicircle is "closed off" by a diameter.

OpenStudy (anonymous):

6300/22?

Directrix (directrix):

@ChiefArnav -->Solve this for r: 90 = pi * r

Directrix (directrix):

How did you get this: 6300/22?

OpenStudy (anonymous):

90 = 22/7 x r r = 90 x 7/22 r = 6300/22 r= 286.36

Directrix (directrix):

Okay, next use the area formula Area semicircle = .5 * pi *r^2 Put your value of r in the formula and crank out the area of the semicircle.

OpenStudy (anonymous):

why .5?

Directrix (directrix):

Because we are finding the area of half of the circle.

Directrix (directrix):

Calculate its area I assume that "its" refers to the semicircle. Do you think "its" refers to the entire circle? Check your problem source.

OpenStudy (anonymous):

then why not pi*r^2/2?

Directrix (directrix):

Same thing. Use that if you like. But do not square the 2 when you square the radius.

Directrix (directrix):

Post your answer so we can compare.

OpenStudy (anonymous):

22/7*28636/100*28636/100?

OpenStudy (anonymous):

all over 2

Directrix (directrix):

What does that crank out to be?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i got 5154111

Directrix (directrix):

That is not what I got.

OpenStudy (anonymous):

can you TYPE out your work?

Directrix (directrix):

No. And, if I did and it looked like this: 22/7*28636/100*28636/100 then it would still make no sense to me.

OpenStudy (anonymous):

i'm confused:(

Directrix (directrix):

What did you get for the radius of the semicircle?

OpenStudy (anonymous):

286.36cm

Directrix (directrix):

I got r = 28.648 for the radius

OpenStudy (anonymous):

oh,no worries i got the right ans:)

Directrix (directrix):

That is what this gives: 90 = 22/7 * r so you should get the same thing for r.

Directrix (directrix):

For the area of the semicircle, you used this: p i * ( (r^2) / 2 )

Directrix (directrix):

Which gives the same answer I got: http://www.wolframalpha.com/input/?i=pi+*+%28+%2828.648%5E2%29+%2F+2+%29 @ChiefArnav